find the equation of the line containing (2,-3) and having the same slope as the line 3x + 4y = 10
Answer by mananth(16946)

3 x + 4 y = 10
Find the slope of this line
make y the subject
4 y = -3 x + 10
Divide by 4
y = -0.75 x + 2.5
Compare this equation with y=mx+b
slope m = - 3/4
The slope of a line parallel to the above line will be the same
The slope of the required line will be -0.75
m= - 3/4 ,point ( 2 , -3 )
Find b by plugging the values of m & the point in
y=mx+b
-3 = -1.50 + b
b= -1.5
m= -0.75
Plug value of the slope and b in y = mx +b
The required equation isy = - 3/4 x -1.5
y = - 3/ 4 x -3/ 2
m.ananth@hotmail.ca